Exercise 5-18 Allocating product cost between cost of goods sold and ending inventory: Multiple purchases LO 5-6 Cortez Company sells chairs that are used at computer stations. Its beginning inventory of chairs was 180 units at $43 per unit. During the year, Cortez made two batch purchases of this chair. The first was a 260-unit purchase at $48 per unit; the second was a 390-unit purchase at $50 per unit. During the period, it sold 490 chairs. Required Determine the amount of product costs that would be allocated to cost of goods sold and ending inventory, assuming that Cortez uses FIFO. LIFO. Weighted average. (Do not round intermediate calculations. Round your final answers to nearest whole dollar amount.)

Required
Determine the amount of product costs that would be allocated to cost of goods sold and ending inventory, assuming that Cortez uses

  1. FIFO.
  2. LIFO.
  3. Weighted average. (Do not round intermediate calculations. Round your final answers to nearest whole dollar amount.)

Find

FIFO LIFO Weighted Average
Cost of goods $ $ $
Ending Inventory $ $ $
Answer #1

Solution:

FIFO

LIFO

Weighted average

Cost of goods sold

22,720

24,300

23,449

Ending Inventory

17,000

15,420

16,271

Working:

Units

Per unit price

Beginning

180

43

First Purchase

260

48

Second Purchase

390

50

FIFO: COGS

Units

Per unit price

Total

180

43

7,740

260

48

12,480

50

50

2,500

490

22,720

FIFO: Ending inventory

Units

Per unit price

Total

340

50

17,000

LIFO: COGS

Units

Per unit price

Total

390

50

19,500

100

48

4,800

490

24,300

LIFO: Ending inventory

Units

Per unit price

Total

180

43

7,740

160

48

7,680

15,420

Weighted average: COGS

Units

Per unit price

Total

Beginning

180

43

7,740

First Purchase

260

48

12,480

Second Purchase

390

50

19,500

830

39,720

Per unit price = 39720 / 830 = 47.86

COGS = 490 * 47.86 = 23,449.16

Ending inventory: 340 * 47.86 = 16,270.84